# Knowledge Independence Breeds Disruption but Limits Recognition

### v1

@@ Acknowledgments
Table S11: Normalized mediation regression on paper impact in SciSciNet dataset. We perform normalized mediation analysis to assess the effect of paper impact on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C5$, team size, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by paper impact, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Citation percentile KI -0.0873*** -0.0188*** -0.1061***
(0.0002) (0.0004) (0.0005)
Citation percentile Novelty -0.0019*** -0.1042*** -0.1061***
(0.0000) (0.0005) (0.0005)
Citation percentile Reference count -0.0456*** -0.0605*** -0.1061***
(0.0002) (0.0005) (0.0005)
Citation percentile Reference age -0.0036*** -0.1024*** -0.1061***
(0.0001) (0.0005) (0.0005)
Citation percentile Reference $C_{5}$0.0004*** -0.1057 -0.1061***
(0.0000) (0.0005) (0.0005)
Citation percentile Team size -0.0009*** -0.1052*** -0.1061***
(0.0002) (0.0005) (0.0005)
Citation percentile Team distance -0.0029*** -0.1032*** -0.1061***
(0.0000) (0.0000) (0.0005)
Citation percentile Team freshness -0.0038*** -0.1023*** -0.1061***
(0.0000) (0.0005) (0.0005)

### v2

@@ Acknowledgements
    We perform normalized confounding analysis to assess the effect of paper impact on disruption in Web of Science dataset. The candidate confounding variables include: KI, novelty, interdisciplinarity, reference count, reference age, team size, team distance, and team freshness, with discipline and publication year controlled. The relative change in estimate (RCE) refers to the relative change of effect after controlling for the confounder. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
S8 Extended analyses to OECD patent dataset
We have replicated our core analyses using the OECD-Patent data [^2] https://www.oecd.org/en/data/datasets/intellectual-property-statistics.html, with the latest data version of Feb-2025. This dataset, known for its extensive coverage and larger volume compared to other patent databases like PatentView, allows for a more comprehensive analysis of global technological innovation trends.
Our analysis of the patent data revealed a distinctive citation landscape. We found that the distribution of KI in patents is highly skewed, with over half of the patents having a maximum KI score of 1.0 (Fig. S27A). The remaining KI scores are also systematically biased towards positive values. This unique distribution reflects a key difference between scientific and technological knowledge systems, which we attribute to the inherent review mechanisms of patent applications. Patent examiners systematically assess the relevance of cited prior art; patents with overly similar or dependent citations are often rejected due to a lack of novelty and inventiveness, which, by design, leads to a systemic preference for citation independence.
We also analyzed the temporal evolution of patent KI, which supports the influence of this institutional mechanism (Fig. S27B). From 1978 to 2010, the average KI of patents showed a clear downward trend, consistent with the decline cycle of the patent disruption displayed in Figure 2 by Park et al. (2023)  [park2023papers]. However, this decline slowed around 2000, coinciding with the initial internationalization of patent examination driven by the cooperation between the European Patent Office (EPO) and the U.S. Patent and Trademark Office (USPTO). Notably, with the formal launch of the Cooperative Patent Classification (CPC) system by the EPO and the USPTO in 2013333European Patent Office, “Highlights 2012: Cooperative Patent Classification”, EPO Annual Report 2012., which signaled a greater convergence of global patent examination standards, KI has since shown an overall upward trend.
Figure S27: $|$ The distribution and evolution pattern of knowledge independence in the OECD patent dataset. (A) The KI distribution of 16,966,216 patents with at least two references published between 1978 and 2024 in the OECD. (B) Patents’ KI exhibit a pervasive downtrend from 1978 to 2010 and upward trend since around 2010 across technical domains (the domain category mapping can be found via https://www.uspto.gov/web/patents/classification/cpc/html/cpc.html). Bootstrapped 95% confidence intervals are shown as shaded bands.
Despite these institutional differences, we found compelling evidence that KI’s relationship with technological innovation mirrors its role in scientific disruption. Specifically, our analysis confirms a significant positive association between KI and patent disruption (Fig. S28A). This finding is robust across various statistical controls (Fig. S28B-L) and is further supported by causal inference analyses (Fig. S29, Table S16). Given the highly skewed distribution of KI, with approximately half of the observations having a value of 1, we conducted a robustness check by binarizing the variable. We created a binary variable representing “fully independence” ($\text{KI}=1$) versus “non-fully independence” ($\text{KI}<1$). Despite this simplification, our analysis continued to find a robust positive association between patent KI and disruption (Fig. S30-S31, Table S17), suggesting that our core finding is not sensitive to this distributional property. These results consistently show that a higher KI, regardless of the institutional context, is a key driver of technological disruption. The consistency of these results across the scientific and technological domains strongly reinforces the generalizability of our conceptual framework and the predictive power of the KI metric on disruption.
Regarding the relationship between KI and patent impact, though it’s a compelling area, we have opted to defer its comprehensive analysis to future work. This decision is based on two key considerations: first, patent impact is arguably more closely linked to commercial value than to citation counts; and second, the central focus of our current study is the citation-based recognition system. A systematic investigation into the connection between KI and patent impact will be a crucial component of our future research agenda.
Figure S28: $|$ Knowledge independence is associated with technical disruption in the OECD patent dataset. (A) For the 11,349,527 patents published before 2023 in the OECD patent dataset that are cited at least once, the average disruption percentile (blue curve, left y-axis) and the disruption positive ratio (green curve, right y-axis) increase with the KI percentile. The association between KI and disruption persists, regardless of the impact $C_{5}$ (panel B), publication decade (panel C), technical domain (panel D), reference count (panel E), reference age (panel F), reference impact $C_{5}$ (panel G), reference disruption (panel H), interdisciplinarity (panel I), team size (panel J), geographic distance (panel K), and collaboration freshness (panel L). Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure S29: $|$ Causal inference by CEM and PSM in the OECD patent dataset. The ATT matrixs of KI on disruption via CEM (A) and PSM (B). Each controlled group is set as a baseline, and ATTs are calculated for comparisons between the baseline and each of the treated groups. Blue cells represent negative ATTs, while red ones represent positive ATTs, with color intensity proportional to the absolute value. Each ATT is tested against the null hypothesis (ATT equals 0) using a two-sided t-test. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Table S16: $|$ Normalized regression models of disruption and the influencing factors in the OECD dataset.
Model 1 Model 2 Model 3 Model 4 Model 5 Model 6 Model 7 Model 8 $R^{2}$%
KI 0.1712*** 0.1711*** 0.1644*** 0.1713*** 0.1707*** 0.1705*** 0.1713*** 0.1638*** 40.65
(0.001) (0.001) (0.001) (0.001) (0.001) (0.001) (0.001) (0.001)
Interdisciplinarity -0.0045*** -0.0028** 0.07
(0.001) (0.001)
Reference count -0.1021*** -0.1012*** 19.52
(0.001) (0.001)
Reference age -0.0092*** 0.0025* 0.76
(0.001) (0.001)
Team size -0.0222*** -0.0037** 0.54
(0.001) (0.001)
Team distance -0.0276*** -0.0215*** 0.74
(0.001) (0.001)
Team freshness -0.0129*** -0.0101*** 0.10
(0.001) (0.001)
Discipline fixed effects Yes Yes Yes Yes Yes Yes Yes Yes 28.82
Year fixed effects Yes Yes Yes Yes Yes Yes Yes Yes 8.79
N 927551 927551 927551 927551 927551 927551 927551 927551
$R^{2}$ 0.065 0.065 0.075 0.065 0.066 0.066 0.065 0.076
  • $\dagger$
    We perform normalized ordinary-least-squares (OLS) regression analyses to examine the effect of KI on disruption in the OECD patent dataset, while controlling for various covariates. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Figure S30: $|$ Binary knowledge independence is associated with technical disruption in the OECD patent dataset. (A) For the 11,349,527 patents published before 2023 in the OECD patent dataset that are cited at least once, the average disruption percentile (blue curve, left y-axis) and the disruption positive ratio (green curve, right y-axis) increase with the KI percentile. The association between KI and disruption persists, regardless of the impact $C_{5}$ (panel B), publication decade (panel C), technical domain (panel D), reference count (panel E), reference age (panel F), reference impact $C_{5}$ (panel G), reference disruption (panel H), interdisciplinarity (panel I), team size (panel J), geographic distance (panel K), and collaboration freshness (panel L). Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure S31: $|$ Causal inference by CEM and PSM in the OECD patent dataset. The ATT matrixs of binary KI on disruption via CEM (A) and PSM (B). Each controlled group is set as a baseline, and ATTs are calculated for comparisons between the baseline and each of the treated groups. Blue cells represent negative ATTs, while red ones represent positive ATTs, with color intensity proportional to the absolute value. Each ATT is tested against the null hypothesis (ATT equals 0) using a two-sided t-test. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Table S17: $|$ Normalized regression models of disruption and the influencing factors in the OECD dataset.
Model 1 Model 2 Model 3 Model 4 Model 5 Model 6 Model 7 Model 8 $R^{2}$%
KI 0.1884*** 0.1885*** 0.1742*** 0.1884*** 0.1877*** 0.1874*** 0.1883*** 0.1739*** 46.59
(0.001) (0.001) (0.001) (0.001) (0.001) (0.001) (0.001) (0.001)
Interdisciplinarity 0.0027* 0.0032** 0.07
(0.001) (0.001)
Reference count -0.0831*** -0.0831*** 16.40
(0.001) (0.001)
Reference age 0.0017 0.0106*** 0.64
(0.001) (0.001)
Team size -0.0198*** -0.0037** 0.50
(0.001) (0.001)
Team distance -0.0244*** -0.0201*** 0.69
(0.001) (0.001)
Team freshness -0.0098*** -0.0081*** 0.08
(0.001) (0.001)
Discipline fixed effects Yes Yes Yes Yes Yes Yes Yes Yes 27.25
Year fixed effects Yes Yes Yes Yes Yes Yes Yes Yes 7.78
N 927551 927551 927551 927551 927551 927551 927551 927551
$R^{2}$ 0.071 0.071 0.077 0.071 0.071 0.071 0.071 0.078
  • $\dagger$
    We perform normalized ordinary-least-squares (OLS) regression analyses to examine the effect of binary KI on disruption in the OECD patent dataset, while controlling for various covariates. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).